# Segment Any Change

### v1

@@ Abstract
Visual foundation models have achieved remarkable results in zero-shot image classification and segmentation, but zero-shot change detection remains an open problem. In this paper, we propose the segment any change models (AnyChange), a new type of change detection model that supports zero-shot prediction and generalization on unseen change types and data distributions. AnyChange is built on the segment anything model (SAM) via our training-free adaptation method, bitemporal latent matching. By revealing and exploiting intra-image and inter-image semantic similarities in SAM’s latent space, bitemporal latent matching endows SAM with zero-shot change detection capabilities in a training-free way. We also propose a point query mechanism to enable AnyChange’s zero-shot object-centric change detection capability. We perform extensive experiments to confirm the effectiveness of AnyChange for zero-shot change detection. AnyChange sets a new record on the SECOND benchmark for unsupervised change detection, exceeding the previous SOTA by up to 4.4% F score, and achieving comparable accuracy with negligible manual annotations (1 pixel per image) for supervised change detection.
@@ 1 Introduction
We demonstrate the zero-shot prediction capabilities of AnyChange on several change detection datasets, including LEVIR-CD (Chen & Shi, 2020), S2Looking (Shen et al., 2021), xView2 (Gupta et al., 2019), and SECOND (Yang et al., 2021). Due to the absence of published algorithms for zero-shot change detection, we also build baselines from the perspectives of zero-shot change proposal, zero-shot object-centric change detection, and unsupervised change detection. AnyChange outperforms other zero-shot baselines implemented by DINOv2 (Oquab et al., 2023) and SAM with different matching methods in terms of zero-shot change proposal and detection. From the unsupervised change detection perspective, AnyChange beats the previous state-of-the-art model, I3PE (Chen et al., 2023), setting a new record of 48.2% F on SECOND. We show some qualitative results in Fig. 1, demonstrating the zero-shot prediction capabilities of AnyChange on a wide range of application scenarios (i.e., urbanization, disaster damage assessment, de-agriculturalization, deforestation, and natural resource monitoring).
The contributions of this paper are summarized as follows:
@@ 2 Related Work
Segment Anything Model for Change Detection. SAM has been used for change detection via a “parameter-efficient fine-tuning” (PEFT) paradigm (Mangrulkar et al., 2022), such as SAM-CD (Ding et al., 2023) that used Fast-SAM (Zhao et al., 2023) as a frozen visual encoder and fine-tuned adapter networks and the change decoder on change detection datasets in a fully supervised way. This model does not inherit the most two important properties of SAM, i.e., promptability and zero-shot generalization. Fine-tuning in a promptable way with large-scale training change data may be achieve these two properties, however, collecting large-scale bitemporal image pairs with class-agnostic change annotations is non-trivial (Tewkesbury et al., 2015; Zheng et al., 2023), thus no such method exists in the current literature. Our work introduces a new and economic adaptation method for SAM, i.e., training-free adaptation, guaranteeing these two properties with zero additional cost, making zero-shot change detection feasible for the first time.
@@ 3.2 Exploring the Latent Space of SAM
Motivated by the experiments in (Kirillov et al., 2023) on probing the latent space of SAM, we known there are potential semantic similarities between mask embeddings in the same natural image. We further explore the latent space for satellite imagery from both intra-image and inter-image perspectives, thus answering the following two questions:
Q1: Do semantic similarities exist on the same satellite image? Empirically, we find they do. We show this semantic similarity in two ways, i.e., visualizing the first components of principal components analysis (PCA) (Oquab et al., 2023) and probing the latent space (Kirillov et al., 2023), as Fig. 3 shows. Observing the first three PCA components, geospatial objects with the same category have a similar appearance in this low-dimensional subspace. This suggests that this satellite image embedding from SAM encodes the semantics of geospatial objects reasonably well. Furthermore, we do the same latent space probing experiment as Kirillov et al. (2023) did, but on satellite images, and present the results in Fig. 3 (bottom). We compute the mask embedding of object proposals and manually select three proposals with different categories (water, building, and bareland) as queries. By matching query embedding with other mask embeddings, we obtain the most similar object proposals with the query proposal. We find that the most similar object proposals mostly belong to the same category as the query.
Q2: Do semantic similarities exist on satellite images of the same location collected at different times? Empirically, we find they do. To verify this, we introduce a new satellite image from the same geographic area but at a different time . The above satellite image is captured at time . Different from the above latent space probing experiment, we use three object proposals with different spatial positions from the image at , as queries. These three queries have the same category, i.e., building. By matching with all proposals from the image at , we obtain three basically consistent results (F of 68.1%0.67% and recall of 96.2%0.66%), as shown in Fig. 4. This suggests that this semantic similarity exists on the satellite images at different times, even though the images have different imaging conditions because they were taken at different times.
@@ 3.3 Bitemporal Latent Matching
where the is the angle between the directions of the mask embeddings.
@@
Since we need to guarantee the temporal symmetry (Zheng et al., 2021, 2022) of class-agnostic change, we propose to match the object proposals bidirectionally. To this end, the change confidence score for the change at from to is also computed, where is computed on the image embedding with the same object proposal . Afterwards, we can match object proposals and bidirectionally, and change proposals with their confidence score are obtained in total. We propose to finally obtain change detection predictions by sorting by confidence scores and selecting the top-k elements or by angle thresholding. The pseudo-code of Bitemporal Latent Matching is in Appendix A.
@@ 3.4 Point Query Mechanism
| Method | Backbone | F | Prec. | Rec. | mask AR | F | Prec. | Rec. | mask AR | F | Prec. | Rec. | mask AR | F | Prec. | Rec. | mask AR |
@@
4 Experiments
4.1 Zero-Shot Object Change Proposals
Metrics: Conventional change detection mainly focuses on pixel-level evaluation using F, precision, and recall. More and more real-world applications have started to focus on instance-level change, also called “change parcel”. Based on this requirement and AnyChange’s capability of predicting instance change, we adapt the evaluation protocol of the zero-shot object proposal (Kirillov et al., 2023) for the zero-shot change proposal since they have the same output structure. The metric is mask AR@1000 (Lin et al., 2014). Note that change proposals are class-agnostic, therefore, for the first three building-centric datasets, we cannot obtain accurate F and precision due to incomplete “any change” annotations. These two metrics are only used for reference and to see whether the model is close to the naive baseline (predict all as “change” class, and vice versa). Here we mainly focus on recall for both pixel- and instance-levels.
@@ 4.1 Zero-Shot Object Change Proposals
Baselines: AnyChange is based on SAM, however, there is no SAM or other VFM111Visual foundation model, like SAM, CLIP, and DINOv2, etc.-based zero-shot change detection model that can be used for comparison in the current literature. For a fair comparison, we build three strong baselines based on DINOv2 (Oquab et al., 2023) (a state-of-the-art VFM) or SAM. The simple baseline is CVA (Bruzzone & Prieto, 2000), a model-free unsupervised change detection method based on norm as dissimilarity and thresholding. We build “DINOv2+CVA”, an improved version with DINOv2 using an idea similar to DCVA (Saha et al., 2019). We build “SAM+Mask Match”, which follows the macro framework of AnyChange and replaces the latent match with the mask match that adopts the IoU of masks as the similarity. “SAM+CVA Match” follows the same idea of AnyChange but adopts the negative norm of feature difference as the similarity to compute pixel-level change map via SAM feature-based CVA first. The instance-level voting is then adopted to obtain change proposals. We also build each “Oracle” version of AnyChange as an upper bound, where we fine-tune SAM with LoRA () (Hu et al., 2022) and train a change confidence score network on each dataset. More implementation details can be seen in Appendix B.1.
@@
Results: We compare the change recall of AnyChange with other zero-shot baselines in Table 1. For pixel-level change recall, AnyChange achieves better recalls than the other two SAM baselines, especially when using a small ViT-B. This is because bitemporal latent matching better measures semantic similarity, i.e., using the angle between two embeddings. We can observe that the gap between AnyChange and SAM baselines gradually reduces as the backbone becomes larger since visual representation capabilities generally become stronger. AnyChange still has better average performance on four datasets, although a stronger representation can close the performance gap to some extent. This highlights the importance of finding the essential semantic difference. Besides, our AnyChange with ViT-H (636M parameters) achieves comparable recalls to CVA with DINOv2 (ViT-G, 1,100M parameters) on four datasets in fewer parameters. On the SECOND dataset, all variants of AnyChange achieve better zero-shot change detection performance than the strong baseline, DINOv2+CVA, and the margin is up to 3.2% F. For instance-level change recall, AnyChange outperforms the other two SAM baselines by a significant margin. This further confirms the effectiveness and superiority of bitemporal latent matching. We observe that the Oracles obtained via supervised learning have superior precision since they learn semantics and dataset bias explicitly, however, this comes with the cost of recall on pixel or instance levels.
@@ 4.2 Zero-shot Object-centric Change Detection
Results: In Fig. 5, we demonstrate the effect of the point query. Without the point query, AnyChange yields class-agnostic change masks, including building changes, vegetation changes, etc. With a single-point query on a building, we can observe that change masks unrelated to the building are filtered out. Further clicking two more buildings to improve the stability of mask embedding, we find the building changes previously missed are successfully recovered. Table 2 quantitatively reflects this mechanism. With a single-point query, the zero-shot performances on three datasets significantly gain F score. This improvement hurts recall as a cost, however it achieves a better trade-off between precision and recall.
@@
| Method | F | Prec. | Rec. | F | Prec. | Rec. | F | Prec. | Rec. |
@@
After increasing to three-point queries, the recalls of the model on three datasets get back to some extent and the model has comparable precision with the single-point query. The zero-shot performances on three datasets gain F further. These results confirm the effectiveness of the point query as a plugin for AnyChange to provide an interactive mode.
@@
| DINOv2+CVA (zero-shot, our impl.) | ViT-G/14 | 41.4 | 26.9 | 35.2 |
@@
4.3 AnyChange as Change Data Engine
Training recipe: On S2Looking, we train the model on its training set with pure pseudo-labels of AnyChange with ViT-H with a single-point query. All training details follow (Zheng et al., 2023) except the loss function. On SECOND, we train the model on its training set with pure pseudo-labels of AnyChange with ViT-H via fully automatic mode, and other details follow (Zheng et al., 2022). For both these two cases, the loss function is BCE loss with a label smoothing of 0.4. See Appendix B.2 for more details.
Supervised Object Change Detection Results. Table 3 presents standard benchmark results on the S2Looking dataset, which is the one of most challenging building change detection datasets. From a data-centric perspective, we use the same architecture with different fine-tuned data. The model architecture adopts ResNet18-based ChangeStar (196) (Zheng et al., 2021) due to its simplicity and good performance. We set the fine-tuning on 100% ground truth as the upper bound, which achieved 66.3% F. We can observe that fin-tuning on the pseudo-labels of AnyChange with ViT-H yields 40.2% F with 3,500-pixel annotations222The training set of S2Looking has 3,500 image pairs. We apply AnyChange with a single-point query for each pair to produce pseudo-labels. Therefore, the number of manual annotations is 3,500 pixels.. Leveraging AnyChange, we achieve 61% of the upper bound at a negligible cost (% full annotations). We also compare it to the model trained with fewer annotations (1% and 0.1%). We find that their performances are reduced by a significant amount, and are inferior to the model trained with pseudo-labels from AnyChange. This confirms the potential of AnyChange as a change data engine for supervised object change detection.
Unsupervised Change Detection Results. AnyChange’s class-agnostic change masks are natural pseudo-labels for unsupervised change detection. We also compare our AnyChange with unsupervised change detection methods. In Table 4, we find that AnyChange with ViT-B in a zero-shot setting improves over the previous state-of-the-art method, I3PE (Chen et al., 2023). To learn the biases on the SECOND dataset, we trained a ChangeStar (1256) model with pseudo-labels of AnyChange on the SECOND training set. The setting follows I3PE (Chen et al., 2023), thus we align their backbone and use ResNet50 for ChangeStar (1256). The results show that two variants of unsupervised ChangeStar (1256) outperform I3PE. Notably, based on pseudo-labels of AnyChange with ViT-B, our model set a new record of 48.2% F on the SECOND dataset for unsupervised change detection. Besides, we obtain some useful insights for unsupervised change detection: (i) deep features from VFMs significantly assist unsupervised change detection models since DINOv2+CVA beats all advanced competitors except I3PE. Before our strong baseline, DINOv2+CVA, CVA has been always regarded as a simple and ineffective baseline for unsupervised change detection. (ii) dataset biases are helpful for in-domain unsupervised change detection since our model trained with pseudo-labels achieves higher performance than these pseudo-labels. This indicates the model learns some biases on the SECOND dataset, which may be change types and style.

### v4

@@ Abstract
Visual foundation models have achieved remarkable results in zero-shot image classification and segmentation, but zero-shot change detection remains an open problem. In this paper, we propose the segment any change models (AnyChange), a new type of change detection model that supports zero-shot prediction and generalization on unseen change types and data distributions. AnyChange is built on the segment anything model (SAM) via our training-free adaptation method, bitemporal latent matching. By revealing and exploiting intra-image and inter-image semantic similarities in SAM’s latent space, bitemporal latent matching endows SAM with zero-shot change detection capabilities in a training-free way. We also propose a point query mechanism to enable AnyChange’s zero-shot object-centric change detection capability. We perform extensive experiments to confirm the effectiveness of AnyChange for zero-shot change detection. AnyChange sets a new record on the SECOND benchmark for unsupervised change detection, exceeding the previous SOTA by up to 4.4% F1 score, and achieving comparable accuracy with negligible manual annotations (1 pixel per image) for supervised change detection. Code is available at https://github.com/Z-Zheng/pytorch-change-models.
@@ 1 Introduction
We demonstrate the zero-shot prediction capabilities of AnyChange on several change detection datasets, including LEVIR-CD (Chen & Shi, 2020), S2Looking (Shen et al., 2021), xView2 (Gupta et al., 2019), and SECOND (Yang et al., 2021). Due to the absence of published algorithms for zero-shot change detection, we also build baselines from the perspectives of zero-shot change proposal, zero-shot object-centric change detection, and unsupervised change detection. AnyChange outperforms other zero-shot baselines implemented by DINOv2 (Oquab et al., 2023) and SAM with different matching methods in terms of zero-shot change proposal and detection. From the unsupervised change detection perspective, AnyChange beats the previous state-of-the-art model, I3PE (Chen et al., 2023), setting a new record of 48.2% F1 on SECOND. We show some qualitative results in Fig. 1, demonstrating the zero-shot prediction capabilities of AnyChange on a wide range of application scenarios (i.e., urbanization, disaster damage assessment, de-agriculturalization, deforestation, and natural resource monitoring). The contributions of this paper are summarized as follows:
@@ 2 Related Work
Segment Anything Model for Change Detection. SAM has been used for change detection via a “parameter-efficient fine-tuning” (PEFT) paradigm (Mangrulkar et al., 2022), such as SAM-CD (Ding et al., 2023) that used Fast-SAM (Zhao et al., 2023) as a frozen visual encoder and fine-tuned adapter networks and the change decoder on change detection datasets in a fully supervised way. This model does not inherit the most two important properties of SAM, i.e., promptability and zero-shot generalization. Fine-tuning in a promptable way with large-scale training change data may achieve these two properties, however, collecting large-scale bitemporal image pairs with class-agnostic change annotations is non-trivial (Tewkesbury et al., 2015; Zheng et al., 2023), thus no such method exists in the current literature. Our work introduces a new and economic adaptation method for SAM, i.e., training-free adaptation, guaranteeing these two properties with zero additional cost, making zero-shot change detection feasible for the first time.
@@ 3.2 Exploring the Latent Space of SAM
Motivated by the experiments in Kirillov et al. (2023) on probing the latent space of SAM, we known there are potential semantic similarities between mask embeddings in the same natural image. We further explore the latent space for satellite imagery from both intra-image and inter-image perspectives, thus answering the following two questions:
Q1: Do semantic similarities exist on the same satellite image? Empirically, we find they do. We show this semantic similarity in two ways, i.e., visualizing the first components of principal components analysis (PCA) (Oquab et al., 2023) and probing the latent space (Kirillov et al., 2023), as Fig. 3 (a) shows. Observing the first three PCA components, geospatial objects with the same category have a similar appearance in this low-dimensional subspace. This suggests that this satellite image embedding from SAM encodes the semantics of geospatial objects reasonably well. Furthermore, we do the same latent space probing experiment as Kirillov et al. (2023) did, but on satellite images, and present the results in Fig. 3 (a) (bottom). We compute the mask embedding of object proposals and manually select three proposals with different categories (water, building, and bareland) as queries. By matching the query embedding with other mask embeddings, we obtain the most similar object proposals with the query proposal. We find that the most similar object proposals mostly belong to the same category as the query.
Q2: Do semantic similarities exist on satellite images of the same location collected at different times? Empirically, we find they do. To verify this, we introduce a new satellite image from the same geographic area but at a different time . The above satellite image is captured at time . Different from the above latent space probing experiment, we use three object proposals with different spatial positions from the image at , as queries. These three queries have the same category, i.e., building. By matching with all proposals from the image at , we obtain three basically consistent results (F1 of 68.1%0.67% and recall of 96.2%0.66%), as shown in Fig. 3 (b). This suggests that this semantic similarity exists on the satellite images at different times, even though the images have different imaging conditions because they were taken at different times.
@@ 3.3 Bitemporal Latent Matching
Since we need to guarantee the temporal symmetry (Zheng et al., 2021, 2022) of class-agnostic change, we propose to match the object proposals bidirectionally. To this end, the change confidence score for the change at from to is also computed, where is computed on the image embedding with the same object proposal . Afterwards, we can match object proposals and bidirectionally, and change proposals with their confidence score are obtained in total. We propose to finally obtain change detection predictions by sorting by confidence scores and selecting the top-k elements or by angle thresholding. The pseudo-code of Bitemporal Latent Matching is in Appendix B.
@@ 3.4 Point Query Mechanism
4 Experiments
4.1 Zero-Shot Object Change Proposals
Metrics: Conventional change detection mainly focuses on pixel-level evaluation using F1, precision, and recall. More and more real-world applications have started to focus on instance-level change, also called “change parcel”. Based on this requirement and AnyChange’s capability of predicting instance change, we adapt the evaluation protocol of the zero-shot object proposal (Kirillov et al., 2023) for the zero-shot change proposal since they have the same output structure. The metric is mask AR@1000 (Lin et al., 2014). Note that change proposals are class-agnostic, therefore, for the first three building-centric datasets, we cannot obtain accurate F1 and precision due to incomplete “any change” annotations. These two metrics are only used for reference and to see whether the model is close to the naive baseline (predict all as “change” class, and vice versa). Here we mainly focus on recall for both pixel- and instance-levels.
@@ 4.1 Zero-Shot Object Change Proposals
| Method | Backbone | F1 | Prec. | Rec. | mask AR | F1 | Prec. | Rec. | mask AR | F1 | Prec. | Rec. | mask AR | F | Prec. | Rec. | mask AR |
@@
Baselines: AnyChange is based on SAM, however, there is no SAM or other VFM111Visual foundation model, like SAM, CLIP, and DINOv2, etc.-based zero-shot change detection model that can be used for comparison in the current literature. For a fair comparison, we build three strong baselines based on DINOv2 (Oquab et al., 2023) (a state-of-the-art VFM) or SAM. The simple baseline is CVA (Bruzzone & Prieto, 2000), a model-free unsupervised change detection method based on norm as dissimilarity and thresholding. We build “DINOv2+CVA”, an improved version with DINOv2 using an idea similar to DCVA (Saha et al., 2019). We build “SAM+Mask Match”, which follows the macro framework of AnyChange and replaces the latent match with the mask match that adopts the IoU of masks as the similarity. “SAM+CVA Match” follows the same idea of AnyChange but adopts the negative norm of feature difference as the similarity to compute pixel-level change map via SAM feature-based CVA first. The instance-level voting is then adopted to obtain change proposals. We also build each “Oracle” version of AnyChange as an upper bound, where we fine-tune SAM with LoRA () (Hu et al., 2022) and train a change confidence score network on each dataset. More implementation details can be seen in Appendix C.1.
@@
Results: We compare the change recall of AnyChange with other zero-shot baselines in Table 1. For pixel-level change recall, AnyChange achieves better recalls than the other two SAM baselines, especially when using a small ViT-B. This is because bitemporal latent matching better measures semantic similarity, i.e., using the angle between two embeddings. We can observe that the gap between AnyChange and SAM baselines gradually reduces as the backbone becomes larger since visual representation capabilities generally become stronger. AnyChange still has better average performance on four datasets, although a stronger representation can close the performance gap to some extent. This highlights the importance of finding the essential semantic difference. Besides, our AnyChange with ViT-H (636M parameters) achieves comparable recalls to CVA with DINOv2 (ViT-G, 1,100M parameters) on four datasets in fewer parameters. On the SECOND dataset, all variants of AnyChange achieve better zero-shot change detection performance than the strong baseline, DINOv2+CVA, and the margin is up to 3.2% F1. For instance-level change recall, AnyChange outperforms the other two SAM baselines by a significant margin. This further confirms the effectiveness and superiority of bitemporal latent matching. We observe that the Oracles obtained via supervised learning have superior precision since they learn semantics and dataset bias explicitly, however, this comes with the cost of recall on pixel or instance levels.
@@
Ablation: Matching Direction. As presented in Table 2, we can find that the performance of single-directional matching is sensitive to temporal order, e.g., mask AR of two single-directional matching on LEVIR-CD are 1.3% and 35.9%, respectively. This is because the class-agnostic change is naturally temporal symmetric which is exactly the motivation of our bidirectional design. This result also confirms generally higher and more robust zero-shot change proposal capability.
Ablation: Robustness to radiation variation. We used ViT-B as the backbone for fast experiments. The results are presented in Table 3. The performance jitter of mask AR is less than 2% (-1.9%,+0.1%, -1.9%, -1.4%) on these four datasets. We believe this sensitivity to radiation variation is acceptable for most applications.
| LEVIR-CD | S2Looking (binary) | xView2 (binary) | SECOND (binary) | |||||||||||||
| Direction | F1 | Prec. | Rec. | mask AR | F1 | Prec. | Rec. | mask AR | F1 | Prec. | Rec. | mask AR | F | Prec. | Rec. | mask AR |
| bidirectional | 23.4 | 13.7 | 83.0 | 32.6 | 7.4 | 3.9 | 94.0 | 48.3 | 13.4 | 7.6 | 59.3 | 27.8 | 44.6 | 30.5 | 83.2 | 27.0 |
| from to | 17.7 | 10.1 | 72.8 | 1.3 | 9.0 | 4.8 | 85.6 | 32.1 | 15.3 | 9.0 | 49.8 | 27.3 | 41.2 | 31.2 | 60.6 | 14.8 |
| from to | 23.6 | 13.6 | 88.7 | 35.9 | 8.1 | 4.3 | 79.5 | 19.4 | 12.3 | 7.6 | 32.7 | 6.7 | 46.1 | 34.1 | 71.3 | 14.9 |
| LEVIR-CD | S2Looking (binary) | xView2 (binary) | SECOND (binary) | |||||||||||||
| Condition | F1 | Prec. | Rec. | mask AR | F1 | Prec. | Rec. | mask AR | F1 | Prec. | Rec. | mask AR | F | Prec. | Rec. | mask AR |
| baseline | 23.4 | 13.7 | 83.0 | 32.6 | 7.4 | 3.9 | 94.0 | 48.3 | 13.4 | 7.6 | 59.3 | 27.8 | 44.6 | 30.5 | 83.2 | 27.0 |
| w/ color jitter | 22.6 | 13.1 | 84.2 | 30.7 | 7.4 | 3.8 | 94.1 | 48.4 | 13.5 | 7.7 | 54.7 | 25.9 | 42.4 | 28.7 | 81.7 | 26.4 |
@@ 4.2 Zero-shot Object-centric Change Detection
Results: We demonstrate the effect of the point query in Fig. 4. Without the point query, AnyChange yields class-agnostic change masks, including building changes, vegetation changes, etc. With a single-point query on a building, we can observe that change masks unrelated to the building are filtered out. Further clicking two more buildings to improve the stability of mask embedding, we find the building changes previously missed are successfully recovered. Table 4 quantitatively reflects this mechanism. With a single-point query, the zero-shot performances on three datasets significantly gain F1 score. This improvement hurts recall as a cost, however it achieves a better trade-off between precision and recall.
@@
| Method | F1 | Prec. | Rec. | F1 | Prec. | Rec. | F1 | Prec. | Rec. |
@@
After increasing to three-point queries, the recalls of the model on three datasets get back to some extent, and the model has comparable precision with the single-point query. The zero-shot performances on three datasets gain F1 further. These results confirm the effectiveness of the point query as a plugin for AnyChange to provide an interactive mode.
4.3 AnyChange as Change Data Engine
Training recipe: On S2Looking, we train the model on its training set with pure pseudo-labels of AnyChange with ViT-H with a single-point query. All training details follow Zheng et al. (2023) except the loss function. On SECOND, we train the model on its training set with pure pseudo-labels of AnyChange with ViT-H via fully automatic mode, and other details follow Zheng et al. (2022). For both these two cases, the loss function is BCE loss with a label smoothing of 0.4. See Appendix C.2 for more details.
@@ 4.3 AnyChange as Change Data Engine
Supervised Object Change Detection Results. Table 5 presents standard benchmark results on the S2Looking dataset, which is the one of most challenging building change detection datasets. From a data-centric perspective, we use the same architecture with different fine-tuned data. The model architecture adopts ResNet18-based ChangeStar (196) (Zheng et al., 2021) due to its simplicity and good performance. We set the fine-tuning on 100% ground truth as the upper bound, which achieved 66.3% F1. We can observe that fin-tuning on the pseudo-labels of AnyChange with ViT-H yields 40.2% F1 with 3,500-pixel annotations222The training set of S2Looking has 3,500 image pairs. We apply AnyChange with a single-point query for each pair to produce pseudo-labels. Therefore, the number of manual annotations is 3,500 pixels.. Leveraging AnyChange, we achieve 61% of the upper bound at a negligible cost (% full annotations). We also compare it to the model trained with fewer annotations (1% and 0.1%). We find that their performances are reduced by a significant amount, and are inferior to the model trained with pseudo-labels from AnyChange. This confirms the potential of AnyChange as a change data engine for supervised object change detection.
Unsupervised Change Detection Results. AnyChange’s class-agnostic change masks are natural pseudo-labels for unsupervised change detection. We also compare our AnyChange with unsupervised change detection methods. In Table 6, we find that AnyChange with ViT-B in a zero-shot setting improves over the previous state-of-the-art method, I3PE (Chen et al., 2023). To learn the biases on the SECOND dataset, we trained a ChangeStar (1256) model with pseudo-labels of AnyChange on the SECOND training set. The setting follows I3PE (Chen et al., 2023), thus we align their backbone and use ResNet50 for ChangeStar (1256). The results show that two variants of unsupervised ChangeStar (1256) outperform I3PE.
@@
| DINOv2+CVA (zero-shot, our impl.) | ViT-G/14 | 41.4 | 26.9 | 89.4 |
@@
Notably, based on pseudo-labels of AnyChange with ViT-B, our model set a new record of 48.2% F1 on the SECOND dataset for unsupervised change detection. Besides, we obtain some useful insights for unsupervised change detection: (i) deep features from VFMs significantly assist unsupervised change detection models since DINOv2+CVA beats all advanced competitors except I3PE. Before our strong baseline, DINOv2+CVA, CVA has been always regarded as a simple and ineffective baseline for unsupervised change detection. (ii) dataset biases are helpful for in-domain unsupervised change detection since our model trained with pseudo-labels achieves higher performance than these pseudo-labels. This indicates the model learns some biases on the SECOND dataset, which may be change types and style.